You are running the synthesis of ammonia reaction: 3 H2(g) + N2(g) --> 2 NH3(g) You add 15 g N2 and 9.0 g H2 to your reaction flask. Which of the following statements are true? (More than one option may be correct)

H2 is the limiting reactant because the limiting reactant calculation showed that less NH3 can be produced from H2 as compared to the amount of NH3 possible from the N2.

N2 is the limiting reactant because the limiting reactant calculation showed that less NH3 can be produced from N2 as compared to the amount of NH3 possible from the H2.

It can be assumed that H2 is limiting because there are fewer grams of H2 in the flask.

It can be assumed that N2 is limiting because there are more grams of N2 in the flask.
Answer #1

number of moles of N2 = 15 g / 28 g/mol = 0.536 moles

number of moles of H2 = 9g / 2g/mole = 4.5 moles

number of moles of N2 needed = (1 mole of N2 * 4.5 mole H2) / (3 mole of H2) = 1.5 moles of N2

but we have only 0.536 mol of N2

so N2 is limiting reagent

because the limiting reactant calculation showed that less NH3 can be produced from N2 as compared to the amount of NH3 possible from the H2